# SplatVoxel: History-Aware Novel View Streaming without Temporal Training

### v1

@@ C.1 Static Scene Datasets
DL3DV [37]
For the DL3DV dataset, we set the training and testing resolution to $384\times 216$. For training, we randomly select one image in the scene as the target, and randomly select four of the nearest eight cameras to the target as inputs. We scale the scene such that the cameras fit within a unit cube. For evaluation, we use every eighth image as the target set, and for each target we use the nearest four cameras not in the target set as inputs. We average metrics per scene, and then average over all scenes. In our experiments on the DL3DV test set (Table 4), both our model and GS-LRM are trained using 8 GPUs, with a batch size of 2 per GPU, for a total of 200K iterations. We first train the multi-view transformer backbone for 100K iterations, followed by 100K iterations of joint training with the voxel decoder.
@@
C.2 Dynamic Scene Datasets
Neural3DVideo [26]
We use six sequences from the dataset for evaluation. The center target camera is selected as the evaluation view, while the two second-closest cameras on either side are used as input cameras. We use resolution $320\times 240$ downsampling from their original resolution.
LongVolumetricVideo [75]
For evaluation, we use the Corgi and Bike sequences and select a 4-input-camera setup that covers one target view. We failed to find enough baseline coverage for other sequences in LongVolCap using only 4 input cameras. For both the Corgi and Bike sequences, we use 1500 frames for evaluation.
We use resolution $384\times 216$ for Corgi and $256\times 256$ for the Bike sequences, downsampling from their original resolution.
Appendix D Baselines
@@
We use the original 3D Gaussian Splatting implementation [22], and reduce the number of training iterations from original 30,000 to 1,000 for Neural3DVideo and 1,500 for LongVolumetricVideo. This modification not only accelerates training but also significantly enhances novel view synthesis results (e.g. around 4 PSNR improvement). Given the sparsity of the input data, this adjustment is crucial for preventing overfitting. All other parameters are kept consistent with the original implementation.
@@
We use the open-source 3DGStream [62] implementation to process the multi-view videos. We first use the original 3DGS to optimize on the first frame for 1,500 iterations, warmup the NTC cache, and then optimize on the next frame using the previous 3DGS for 150 iterations, following the original implementation.
@@
We use the open-source 4DGS [73] implementation to process the multi-view videos. Following their implementaion, we first optimized the first frame for 3,000 iterations, and then optimize on the full multi-view video sequence for 14,000 iterations.
@@
We enhance GS-LRM’s speed by replacing the CNN deconvolution with an MLP unpatchify layer as proposed in [20], and implementing it in JAX. We maintain the same network parameters and training loss weights as in the original paper. Additionally, we use the LPIPS loss employed in our model, achieving improved LPIPS and SSIM scores, with a slight decrease in PSNR, as detailed in Table 2. We adopt the same color modeling approach as the official paper for RealEstate10K, setting the spherical harmonics degree to 0. For the GS-LRM baseline trained on DL3DV, we set the degree to 1 to ensure a fair comparison with our method. Additionally, we found that this modification enhances GS-LRM’s performance on DL3DV.
Appendix E Additional Results
@@
Table 7 presents detailed metrics for our results alongside the two best-performing baselines, 4DGS and GS-LRM, evaluated across all eight sequences.
@@
Figure 9: Re10K Static scene reconstruction. Additional qualitative examples of GS-LRM trained on RealEstate10k [82].

### v2

@@ B.1 Static Scene Datasets
DL3DV ling2024dl3dv
For the DL3DV dataset, we set the training and testing resolution to $384\times 216$. For training, we randomly select one image in the scene as the target, and randomly select four of the nearest eight cameras to the target as inputs. We scale the scene such that the cameras fit within a unit cube. For evaluation, we use every eighth image as the target set, and for each target we use the nearest four cameras not in the target set as inputs. We average metrics per scene, and then average over all scenes. In our experiments on the DL3DV test set (Table 2), both our model and GS-LRM are trained using 8 GPUs, with a batch size of 2 per GPU, for a total of 200K iterations. Our method supports both joint training of the multi-view Transformer and voxel decoder from scratch, and a two-stage training scheme. As shown in Table 8, joint training from scratch outperforms the two-stage approach, where the multi-view Transformer is first trained for 100K iterations, followed by another 100K iterations of joint training with the voxel decoder. Figure 7 illustrates that after training the multi-view Transformer used in GS-LRM, our proposed sparse voxel Transformer converges to significantly improved results with only a small number of additional training steps.
Table 13: Comparison with Gaussian Graph Network (GGN) zhang2024gaussianGNN. Both models are trained on RealEstate10K (2-view training, 4-view testing).
[…]
GGN zhang2024gaussianGNN 24.76 0.784 0.172
Ours 28.79 0.914 0.081
Table 14: Extension to DepthSplat. Our method can be seamlessly integrated into other feed-forward 3D Gaussian Splatting frameworks to enhance their performance. Both models are trained on a small subset of DL3DV (2-view training, 4-view testing).
[…]
DepthSplat xu2025depthsplat 16.57 0.4108 0.4272
DepthSplat + Ours 17.66 0.4426 0.4019
Figure 7: Comparison of Multi-View and Voxel Transformers. Validation curves on the DL3DV dataset show that our 3D Sparse Voxel Transformer converges faster and achieves significantly better final performance when initialized with a pre-trained 2D Multi-View Transformer, compared to training with the 2D Multi-View Transformer alone. Our method also supports joint training of the Multi-View and Voxel Transformers, leading to further performance improvements as shown in Table 8.
@@
B.2 Dynamic Scene Datasets
For evaluation, we use two input views with a small-baseline setup and across six sequences from Neural3DVideo, and four input views with a large-baseline setup and 1,500 frames for each of the two sequences from LongVolumetricVideo.
Neural3DVideo li2022neural3DV
We use six sequences from the dataset for evaluation, each containing 290 frames. The center target camera is selected as the evaluation view, while the two second-closest cameras on either side are used as input cameras. We use resolution $320\times 240$ downsampling from their original resolution.
LongVolumetricVideo xu2024longvolcap
For evaluation, we use the Corgi and Bike sequences and select a 4-input-camera setup that covers one target view. We failed to find enough baseline coverage for other sequences in LongVolCap using only 4 input cameras. For both the Corgi and Bike sequences, we use 1500 frames for evaluation. We use resolution $384\times 216$ for Corgi and $256\times 256$ for the Bike sequences, downsampling from their original resolution.
Appendix C Baselines
@@
We use the original 3D Gaussian Splatting implementation kerbl20233dgs, and reduce the number of training iterations from original 30,000 to 1,000 for Neural3DVideo and 1,500 for LongVolumetricVideo. This modification not only accelerates training but also significantly enhances novel view synthesis results (e.g. around 4 PSNR improvement). Given the sparsity of the input data, this adjustment is crucial for preventing overfitting. All other parameters are kept consistent with the original implementation.
@@
We use the open-source 3DGStream sun20243dgstream implementation to process the multi-view videos. We first use the original 3DGS to optimize on the first frame for 1,500 iterations, warm up the NTC cache, and then optimize on the next frame using the previous 3DGS for 150 iterations, following the original implementation.
@@
We use the open-source 4DGS wu20244d implementation to process the multi-view videos. Following their implementation, we first optimize the first frame for 3,000 iterations, and then optimize on the full multi-view video sequence for 14,000 iterations.
@@
We enhance GS-LRM’s speed by replacing the CNN deconvolution with an MLP unpatchify layer as proposed in jin2024lvsm, and implementing it in JAX. We maintain the same network parameters and training loss weights as in the original paper. Additionally, we use the LPIPS loss employed in our model, achieving improved LPIPS and SSIM scores, with a slight decrease in PSNR, as detailed in Table 10. We adopt the same color modeling approach as the official paper for RealEstate10K, setting the spherical harmonics degree to 0. For the GS-LRM baseline trained on DL3DV, we set the degree to 1 to ensure a fair comparison with our method. Additionally, we found that this modification enhances GS-LRM’s performance on DL3DV.
Appendix D Additional Results
@@
Table 11 presents detailed metrics for our results alongside the two best-performing baselines, 4DGS and GS-LRM, evaluated across all eight sequences. We show close-ups of the reconstruction in Fig. 8 comparing ours with GS-LRM.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
GT GS-LRM Ours
Figure 8: Reconstruction Close-ups on Dynamic Scenes. We show zoomed-in novel view reconstructions of GS-LRM and our model on dynamic scene datasets. Our model better handles occlusion boundaries with sharper detail.
@@
Figure 10: Re10K Static scene reconstruction. Additional qualitative examples of GS-LRM trained on RealEstate10k zhou2018rel10k.
We also compare our method with Gaussian Graph Network (GGN) zhang2024gaussianGNN on the RealEstate10K dataset, following its experimental configuration. Both models are trained with 2 input views and evaluated with 4 input views to assess generalization across varying input configurations. As shown in Table 13, our method significantly outperforms this state-of-the-art Gaussian fusion approach.
We further validate the generalization capability of the proposed Fuse-and-Refine module by integrating it into DepthSplat xu2025depthsplat. We retrain the publicly available DepthSplat implementation on the DL3DV dataset at a reduced resolution of $128{\times}192$. The training subset includes 1,000 scenes with 2 input views, and evaluation is performed on 10 test scenes with 4 input views. Table B.1 shows that incorporating our module leads to a performance gain over the original DepthSplat, demonstrating the adaptability of our approach.
Appendix E Societal Impact
In this section, we reflect on the broader societal implications of our work. Our method facilitates fast and high-quality 3D reconstruction from sparse views, significantly lowering the barrier for creating digital assets from real-world scenes or objects. This has potential benefits across domains such as digital content creation, robotics, autonomous navigation, cultural heritage preservation, and immersive applications in AR/VR. Additionally, it supports real-time streaming use cases, toward fully immersive experiences.
However, like many 3D/4D reconstruction methods, our approach could be misused to create detailed models of people, places, or objects without consent, raising concerns about privacy, identity theft, and misleading content. While our work is intended for positive and legitimate use, we acknowledge the ethical responsibilities involved and urge users to apply this technique responsibly.